# Semantic Tree Inference on Text Corpa using a Nested Density Approach together with Large Language Model Embeddings

### v1

@@ Abstract
Semantic text classification has undergone significant advances in recent years due to the rise of large language models (LLMs) and their high dimensional embeddings. While LLM-embeddings are frequently used to store and retrieve text by semantic similarity in vector databases, the global structure semantic relationships in text corpora often remains opaque.
Herein we propose a nested density clustering approach, to infer hierarchical trees of semantically related texts. The method starts by identifying texts of strong semantic similarity as it searches for dense clusters in LLM embedding space. As the density criterion is gradually relaxed, these dense clusters merge into more diffuse clusters, until the whole dataset is represented by a single cluster - the root of the tree. By embedding dense clusters into increasingly diffuse ones, we construct a tree structure that captures hierarchical semantic relationships among texts.
We outline how this approach can be used to classify textual data for abstracts of scientific abstracts as a case study. This enables the data-driven discovery research areas and their subfields without predefined categories. To evaluate the general applicability of the method, we further apply it to established benchmark datasets such as the 20 Newsgroups and IMDB 50k Movie Reviews, demonstrating its robustness across domains.
Finally we discuss possible applications on scientometrics, topic evolution, highlighting how nested density trees can reveal semantic structure and evolution in textual datasets.
1 Background
Text classifications date back at least to bag of words models [bagofwords]. In these early approaches, each text is represented as a vector of word frequencies, disregarding the order of words. However, word order is a key semantic feature as changing it can completely alter the meaning of a phrase.
Subsequent refinements introduced term frequency and inverse document frequency (tf-idf) normalization [tfidf], which weights words by their relative importance within a corpus. Later, Latent Semantic Analysis (LSA) [lsa-first, lsa, lsa-intro] applied singlar value decomposition to capture latent relationships among words and documents, enabling more robust document clustering and topic modeling. Easier to scale and based on probabilistic properties instead of linear algebra Latent Dirichlet Allocation (LDA) [lda] and Hirachical Dirichlet Allocation (HDA) [hda] have overtaken the field of topic modelling. Neural network based approaches such as word2vec [word2vec] have arrived in the middle of the last decade. These models could learn the relationships between words in a semantic proximity sense, but still not in a grammatical sequence. The rise of the transformer model [transformer] and the popularization of LLMs allowed for the use of LLM embeddings for document classification, and document retrieval. The transformer model is superieur as it efficiently encodes and uses the position of the words in the text as it generates embeddings. LLM embeddings have a such a vast field of applications and [llm-embed-chatbot, llm-embed-clust] commonly found in retrieval augmented generation (RAG) pipelines [rag, llm-embed-rag]. All of the aforementioned methods project text into a mathematical embedding space, where pairwise distances and clustering can be applied to reveal latent semantic structures. In parallel, hierarchical density clustering techniques have proven highly effective in fields such as biological species classification, phylogenetics [mnhn-tree-tools], and molecular dynamics, where they are used to analyze protein conformations and structural evolution [md-tree-tools].
Here, we extend these hierarchical density clustering methods to textual embeddings, constructing tree-based representations of semantically related texts. This approach goes beyond flat clustering by revealing how topics, subjects, or emotions (depending on the embedding model) relate to one another within a coherent hierarchical framework.
@@ 2 Introduction
Topic modeling and semantic clustering have long aimed to uncover the latent structure of large text corpora. With the rise of large language models (LLMs) and their embeddings, it has become possible to represent textual semantics with unprecedented precision. However, most embedding-based clustering approaches produce flat partitions of the semantic space, providing limited insight into how topics or concepts are hierarchically related or evolve across scales.
In this work, we introduce a hierarchical density-based clustering approach that infers tree structures from LLM embedding spaces. The method builds upon previous applications of hierarchical density clustering in bioinformatics and molecular dynamics, where it has been successfully used to construct phylogenetic or conformational trees from complex, high-dimensional data [mnhn-tree-tools, md-tree-tools]. By adapting this framework to text embeddings, we enable the reconstruction of semantic hierarchies. We show how dense clusters of highly similar documents merge into broader, more diffuse clusters as the similarity threshold is relaxed.
To construct such hierarchical structures, we employ a nested density-based clustering procedure inspired by the DBSCAN algorithm [dbscan]. The core idea is to identify compact, high-density regions of semantically similar embeddings and then progressively relax the density criterion to reveal broader, more diffuse groupings. This iterative relaxation enables clusters discovered at a high-density level to merge into larger clusters at lower density levels, naturally producing a hierarchical tree. Formally, clusters of semantically similar texts that that are densly connected in their respective embedding space are determined by the condition:
$\rho_{L}>\frac{\mathrm{minpts}}{V(\epsilon_{L})},$ (1)
where $\rho_{L}$ represents the denstiy at a given iteration $L$. The right hand side is given by starting parameters to the well known Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [dbscan], with $\epsilon_{L}$ defining a radius at iteration $L$ and $\mathrm{minpts}$ a minimum number of points within this radius, effectively defining the density of a continious dense ensemble to be found by the algorithm. To explore the semantic structure at multiple scales, the radius $\epsilon_{L}$ is incrementally expanded by a small step $\Delta\epsilon$ in each iteration. This results in a sequence of density thresholds: As $\epsilon$ increases, the volume in equation (1) expands leading to lower densities and the identification of more diffuse clusters in each iteration. These new clusters encapsulate denser ones from previous iterations, allowing us to embed fine-grained semantic clusters into coarser ones, thereby forming a hierarchical tree of semantic relationships. For a subsequent layer (iteration) of the tree $L+1$ it follows, equation (2):
$\rho_{L+1}>\frac{\mathrm{minpts}}{V(\epsilon_{L}+\Delta\epsilon)},$ (2)
or alternatively $\epsilon_{L}=\epsilon_{L-1}+\Delta\epsilon$. In our previous work [mnhn-tree-tools], we demonstrated that using this algorithm allows us to construct a hierarchical tree structure by embedding dense clusters into progressively more diffuse clusters until all elements of the conformational space merge into a single cluster - the root of the tree. Figure 1 outlines the tree building procedure conceptually. On the left, a text embedding space is shown, where each colored sphere represents a piece of text. On the right the tree built by the algorithm, from the text embedding space on the left, is shown. One can observe how the grape, and orange colored texts fusion as the density criteria is relaxed into the blue ensemble, encompassing both clusters. At an intermediate step the green cluster stays separate, while the violet texts are not encompassed and seen as outliers. Finally, as the density criterion is further relaxed all texts are only part of one cluster, represented in gray on the right tree-outline.
Figure 1: Tree building from a dataset represented by its embedding space. On the left, an embedding space with clusters of different densities is shown. One can see that the orange, and grape colored clusters embed into a larger blue ensemble. A conceptual tree built from such a two dimensional embedding space is shown on the right.
Our algorithm allows an organization for instance to show how documents within the organization are related. It initially finds separated clusters of highly semantically similar documents (the outer leafs of the tree), that merge, layer by layer together to more diffuse clusters, until all documents form a single cluster (the root of the tree).
We demonstrate our approach on several corpora, including collections of scientific abstracts and benchmark datasets such as 20 Newsgroups and IMDB 50K reviews. In these experiments, we show how the inferred semantic trees capture both fine-grained clusters of related documents and higher-level thematic organization. Applied to the publication records of institutions such as TU Wien and the American University of Beirut, our method reveals naturally emerging research areas and structural relationships between topics.
Once a semantic tree is constructed, an important challenge lies in interpreting and annotating its clusters. We address this by leveraging LLMs themselves: by prompting a model with the text content of each cluster, we can automatically generate concise, human-readable summaries or titles for each node in the hierarchy. This creates an interpretable and fully automated pipeline for semantic structure discovery in large text corpora.
3 Materials and Methods
@@ 3.1 Datasets
  • 20 Newsgroups: The 20 Newsgroups dataset is a widely used benchmark corpus for text classification and topic modeling. It consists of approximately 20,000 newsgroup posts collected from 20 distinct discussion groups on topics such as science, politics, sports, religion, and technology [20newsgroups].
  • IMDB 50K Reviews: A corpus of 50,000 movie reviews from the Internet Movie Database (IMDB), each labeled with a binary sentiment (positive or negative) [50kmovies].
  • AG News: A large-scale news dataset consisting of short article titles and descriptions categorized into four topics: World, Sports, Business, and Science/Technology. [agnews].
@@
Table 1: Statistics about the text corpa used. n(Texts) is the number of unique texts in the dataset, mean and stddev represents the mean and standard deviation of the number of words contained in each text of the dataset.
@@ 3.2 Embeddings
Embeddings for the datasets outlined in section 3.1 were generated using either the Qwen3-Embedding-8B [qwen3embedding] or SFR-Embedding-Mistral [SFRembedding] model. These models were chosen selected based on their high performance in the MTEB-Leaderboard found on Huggingface, and the their practical deployability on the computing infrastructure accessible to us.
4096 dimensonal embedding vectors obtained for each text from a dataset were stored in an in house developed vector database.
@@ 3.3.2 Adaptive Density Clustering with DBSCAN
For each dataset, we constructed trees using an adaptive density-based clustering procedure built upon successive runs of the DBSCAN algorithm. Each DBSCAN iteration identifies clusters at a specific density level, controlled by a triplet of parameters: the initial neighborhood radius $\epsilon_{0}$, the incremental radius step $\Delta\epsilon$, and the minimum number of points within the radius ($\mathrm{minpts}$).
In each successive run, the radius is increased by $\Delta\epsilon$ while keeping $\mathrm{minpts}$ constant. This progressively lowers the effective density threshold and allows previously dense clusters to merge into more diffuse ones. Whenever the number of clusters found in a given iteration is smaller than in the preceding step, the newly identified clusters are stored, and the clusters from the prior iteration are embedded within them. This iterative embedding process builds the hierarchical tree structure described in Figure 1, where dense clusters at outer layers merge stepwise into broader, semantically diffuse clusters toward the root.
@@ 3.4 Tree Annotation
To generate these annotations, we employed a Large Language Model (LLM) capable of handling large textual contexts. Specifically, we used the Llama-4-Scout-17B-16E-Instruct model [meta2024llama4] in 6-bit quantized form, which provides a nominal context window of up to 11 million tokens. Due to technical constraints, our inference pipeline successfully operated with a maximum context size of approximately 600,000 tokens.
@@
In the case were the cluster had to be sampled because it contained more than 1500 abstracts all resulting labels were subsequently fed into a second LLM pass. Again the gpt-oss-20b model [gpt-oss] in 8-bit quantized form is used in such a case where the model inferred the overarching research field encompassing all $s$ subfields. An overview about the different paths that annotation might undergo is outlined in Figure 2.
Figure 2: Shematic represantation of the tree node / cluster annotation process
Tree annotation is computationally intensive. The AUB tree required annotating 33,862 text files, while the TU Wien tree required 62,012, including oversampling. Inference was performed using the llama.cpp [llama-cpp] engine on the MUSICA supercomputer (Austria). Up to 80 NVIDIA H100 GPUs were employed concurrently during annotation. Each compute node with four H100 GPUs (80 GB VRAM each) handled batches of approximately 600,000 tokens per inference instance. Multi-node inference, which could allow for a higher context token limit, was not attempted as MUSICA is currently in its test phase; instead, parallelism was achieved by distributing text files across multiple nodes for simultaneous inference.

### v2

# Discovering Multi-Scale Semantic Structure in Text Corpora Using Density-Based Trees and LLM Embeddings

@@ Abstract
Recent advances in large language models enable documents to be represented as dense semantic embeddings, supporting similarity-based operations over large text collections. However, many web-scale systems still rely on flat clustering or predefined taxonomies, limiting insight into hierarchical topic relationships. In this paper we operationalize hierarchical density modeling on large language model embeddings in a way not previously explored. Instead of enforcing a fixed taxonomy or single clustering resolution, the method progressively relaxes local density constraints, revealing how compact semantic groups merge into broader thematic regions. The resulting tree encodes multi-scale semantic organization directly from data, making structural relationships between topics explicit. We evaluate the hierarchies on standard text benchmarks, showing that semantic alignment peaks at intermediate density levels and that abrupt transitions correspond to meaningful changes in semantic resolution. Beyond benchmarks, the approach is applied to large institutional and scientific corpora, exposing dominant fields, cross-disciplinary proximities, and emerging thematic clusters. By framing hierarchical structure as an emergent property of density in embedding spaces, this method provides an interpretable, multi-scale representation of semantic structure suitable for large, evolving text collections.
@@ 1 Introduction
Understanding the semantic organization of large text corpora is a central challenge in knowledge representation, information retrieval, and semantic web research. As digital collections continue to grow in size and diversity, there is an increasing need for methods that not only group documents by similarity but also expose how concepts relate to one another across multiple levels of abstraction. Such structure is essential for tasks ranging from exploratory search and corpus navigation to the analysis of disciplinary landscapes and knowledge evolution.
Recent advances in large language models (LLMs) have substantially improved the representation of textual meaning. High-dimensional embedding models encode semantic similarity with remarkable fidelity and have become a standard component of modern information systems, particularly in vector databases and semantic retrieval pipelines. While these embeddings enable effective similarity search and flat clustering, they typically provide limited insight into the global organization of semantic space. In particular, they do not directly reveal how fine-grained concepts combine into broader themes, nor how semantic resolution changes as one moves from local to global structure.
Hierarchical representations offer a natural solution to this limitation. By organizing documents into nested structures, hierarchies can capture relationships between topics at multiple semantic scales. Traditional hierarchical approaches in text analysis, however, often rely on strong modeling assumptions, predefined taxonomies, or global distance matrices that are sensitive to noise and high dimensionality. As a result, they can struggle to scale to large corpora or to adapt flexibly to heterogeneous textual content.
In this work, we propose a data-driven approach for reconstructing semantic hierarchies directly from text embedding spaces. Rather than imposing a predefined ontology or topic structure, we infer in an unsupervised manner hierarchical relationships by progressively revealing structure in the embedding space itself. The core idea is to identify groups of highly similar, documents and to examine how these groups merge as semantic similarity criteria are gradually relaxed. This process yields a tree-like structure in which terminal branches correspond to tightly related (dense) texts, while higher-level nodes represent increasingly (diffuse) general semantic groupings.
A key advantage of this approach is that it produces a single, unified representation that simultaneously captures local semantic coherence and global thematic organization. The resulting trees are not restricted to a fixed number of clusters or levels and naturally accommodate documents that do not belong to any dense semantic group at finer resolutions. Such behavior is particularly important for real-world corpora, where topical boundaries are often fuzzy and unevenly populated.
We demonstrate the utility of this framework on several benchmark datasets commonly used in text classification and topic analysis, including 20 Newsgroups, IMDB 50K reviews, and AG News. These experiments allow us to assess how well the inferred hierarchies align with established labels and to analyze how semantic agreement varies across hierarchical depth. Beyond benchmarks, we apply the method to large collections of scientific abstracts from academic institutions, where no single ground-truth taxonomy is assumed. In these cases, the inferred trees reveal institution-specific research profiles, cross-disciplinary connections, and emerging thematic areas.
To support interpretability, we further integrate large language models as semantic annotators. By summarizing the textual content associated with individual tree nodes, we generate concise, human-readable descriptions of semantic regions within the hierarchy. This results in an end-to-end pipeline that combines embedding-based structure discovery with automated semantic labeling.
Overall, this work contributes a flexible and interpretable method for uncovering hierarchical semantic structure in large text corpora. By emphasizing multi-scale organization, data-driven discovery, and semantic transparency, the proposed approach aligns naturally with the goals of semantic web research and provides new tools for exploring complex textual knowledge spaces.
2 Related Work
2.1 Text Representation for Semantic Analysis
Text classification and organization have long relied on vector-based representations to capture semantic relationships within document collections. Early approaches such as bag-of-words models represent texts as unordered collections of terms, enabling efficient comparison but ignoring syntactic structure and semantic context [11]. Subsequent developments introduced weighting schemes such as term frequency–inverse document frequency (tf–idf), which improved discrimination by accounting for term importance within a corpus [30].
Latent Semantic Analysis (LSA) extended these ideas by projecting document-term matrices into lower-dimensional latent spaces using singular value decomposition, enabling the discovery of latent semantic associations [7, 10, 17]. Probabilistic topic models, most notably Latent Dirichlet Allocation (LDA) and its hierarchical extensions, introduced generative frameworks for uncovering thematic structure in document collections [4, 31]. While effective for identifying coarse-grained topics, these models require predefined assumptions about topic structure and often struggle with interpretability at fine semantic resolution [5].
Neural embedding models, including word2vec and related architectures, marked a transition toward distributed semantic representations learned from large corpora [23]. However, these approaches primarily capture word-level proximity and do not fully encode compositional or contextual meaning. The introduction of transformer-based architectures [32] and large language models (LLMs) enabled the generation of contextualized document-level embeddings that capture syntactic structure, semantics, and discourse-level information.
LLM embeddings are now widely used in information retrieval, semantic search, and retrieval-augmented generation pipelines [18, 6]. Despite their expressive power [9, 3, 2] their use in information organization is often limited to nearest-neighbor retrieval or flat clustering, leaving the global semantic structure of document collections largely unexplored.
2.2 Clustering and Hierarchical Organization of Text
Beyond representation, organizing documents into meaningful structures is a core challenge in information processing. Flat clustering methods such as k-means [20, 14] or DBSCAN [8] are frequently applied to document embeddings to identify groups of related texts. While computationally efficient, flat partitions impose a single resolution of analysis and obscure relationships between topics at different levels of semantic abstraction.
Hierarchical approaches aim to overcome this limitation by explicitly modeling multi-level structure. Agglomerative hierarchical clustering constructs dendrograms from pairwise distances but requires global distance matrices and linkage criteria that are sensitive to noise, scaling, and high dimensionality. Similarly, hierarchical topic models rely on strong probabilistic assumptions and often exhibit limited robustness when applied to large, heterogeneous corpora [15, 25].
Density-based clustering offers an alternative by identifying regions of high local similarity without assuming convex cluster shapes. While such methods have been successfully applied in spatial data analysis and other high-dimensional domains [13, 12], their use in text and information retrieval has largely focused on extracting flat clusterings rather than interpreting the full hierarchical structure. As a result, the potential of density-based hierarchies for revealing multi-scale semantic organization in text corpora remains to be explored.
2.3 From Density-Based Clustering to Semantic Trees
Hierarchical density-based clustering has been successfully applied in other high-dimensional domains, such as bioinformatics and molecular dynamics, to reconstruct phylogenetic relationships and conformational landscapes [13, 12]. In these settings, trees are constructed by progressively relaxing density constraints, embedding compact structures into broader ones and yielding interpretable representations of multi-scale structure.
This work adapts this paradigm to the organization of document embeddings. Rather than treating the hierarchy as an intermediate artifact or selecting a single optimal partition, we construct and analyze the full density-based tree as the primary representation of semantic structure. The resulting semantic tree captures how documents group and merge across similarity scales, enabling analysis at multiple levels of abstraction, from fine-grained topical distinctions to broad thematic domains.
Crucially, the construction is independent of predefined taxonomies or label sets. This makes the approach well suited to exploratory analysis of large and heterogeneous corpora, where existing classifications are often incomplete, inconsistent, or entirely absent.
2.4 Interpretability and Automatic Annotation
Hierarchical organization alone is insufficient without interpretability. Large trees derived from embedding spaces require semantic summaries that allow users to understand and navigate the structure. Recent advances in LLMs enable large-context inference, making it possible to generate concise semantic descriptions for collections of documents.
We leverage this capability to automatically annotate tree nodes by prompting LLMs with the textual content of each cluster. This produces human-readable labels that summarize the dominant themes represented at different levels of the hierarchy. The result is an end-to-end pipeline that combines unsupervised structure discovery with automatic semantic interpretation, addressing a long-standing challenge in large-scale document organization.
3 Methodology
@@ 3.1 Datasets
  • 20 Newsgroups: The 20 Newsgroups dataset is a widely used benchmark corpus for text classification and topic modeling. It consists of approximately 20,000 newsgroup posts collected from 20 distinct discussion groups on topics such as science, politics, sports, religion, and technology [24].
  • IMDB 50K Reviews: A corpus of 50,000 movie reviews from the Internet Movie Database (IMDB), each labeled with a binary sentiment (positive or negative) [19].
  • AG News: A large-scale news dataset consisting of short article titles and descriptions categorized into four topics: World, Sports, Business, and Science/Technology. [33].
@@
Table 1: Statistics about the text corpa used. n(Texts) is the number of unique texts in the dataset, mean and stddev represent the mean and standard deviation of the number of words contained in each text of the dataset.
@@ 3.2 Embeddings
Embeddings for the datasets outlined in section 3.1 were generated using either the Qwen3-Embedding-8B [34] or SFR-Embedding-Mistral [22] model. These models were chosen selected based on their high performance in the MTEB-Leaderboard found on Huggingface, and their practical deployability on the computing infrastructure accessible to us.
4096-dimensonal embedding vectors obtained for each text from a dataset were stored in an in house developed vector database.
@@ 3.3.2 Adaptive Density Clustering with DBSCAN
Semantic clustering of text documents was performed using a density-based method adapted from DBSCAN [8]. At iteration $L$, clusters were defined as regions where the local density $\rho_{L}$ satisfies
$\rho_{L}>\frac{\mathrm{minpts}}{V(\epsilon_{L})},$ (1)
where $\epsilon_{L}$ is the neighborhood radius defining a volume $V(\epsilon_{L})$, and $\mathrm{minpts}$ is the minimum number of points required within that volume. Unlike HDBSCAN [21], which constructs a hierarchy internally but typically outputs a single flat clustering, this method explicitly preserves hierarchical structure across successive density thresholds.
To capture semantic structure at multiple scales, $\epsilon_{L}$ was increased in fixed increments $\Delta\epsilon$, generating progressively lower density thresholds. Clusters identified at each step contained denser clusters from previous iterations. For iteration $L+1$, the density criterion becomes
$\rho_{L+1}>\frac{\mathrm{minpts}}{V(\epsilon_{L}+\Delta\epsilon)}.$ (2)
For each dataset, hierarchical trees were constructed using successive runs of DBSCAN. Each iteration is controlled by three parameters: the initial radius $\epsilon_{0}$, the incremental step $\Delta\epsilon$, and the minimum points $\mathrm{minpts}$. In each successive run, the radius is increased while keeping $\mathrm{minpts}$ constant, progressively lowering the effective density threshold. Previously dense clusters merge into larger, more diffuse clusters, and whenever fewer clusters are found than in the preceding iteration, the new clusters are stored, embedding the previous ones within them. Figure 1 illustrates this process conceptually.
Figure 1: Tree construction from a dataset represented in embedding space. Left: embedding space with clusters of different densities; each colored sphere represents a text document. Right: conceptual tree built from the embedding space, showing how denser clusters (orange and grape) merge into larger ensembles (blue) as the density threshold is relaxed, while other clusters (green) remain separate and some points (violet) are considered outliers. Eventually, all points merge into a single root cluster (gray).
Unlike methods that optimize a predefined objective (e.g., topic coherence, likelihood, or cluster compactness), this approach constructs a hierarchical semantic representation directly from local density relations in the embedding space. It allows semantic groupings to emerge across multiple resolutions without assuming a fixed taxonomy, linkage criterion, or global distance structure. To our knowledge, this is the first method that operationalizes nested density relaxation as a general-purpose mechanism for semantic tree induction in text corpora.
@@ 3.4 Tree Annotation
To generate these annotations, we employed a Large Language Model (LLM) capable of handling large textual contexts. Specifically, we used the Llama-4-Scout-17B-16E-Instruct model [1] in 6-bit quantized form, which provides a nominal context window of up to 11 million tokens. Due to technical constraints, our inference pipeline successfully operated with a maximum context size of approximately 600,000 tokens.
@@
In the case where the cluster had to be sampled because it contained more than 1500 abstracts all resulting labels were subsequently fed into a second LLM pass. Again the gpt-oss-20b model [28] in 8-bit quantized form is used in such a case where the model inferred the overarching research field encompassing all $s$ subfields. An overview about the different paths that annotation might undergo is outlined in Figure 2.
Figure 2: Schematic representation of the tree node / cluster annotation process.
Tree annotation is computationally intensive. The AUB tree required annotating 33,862 text files, while the TU Wien tree required 62,012, including oversampling. Inference was performed using the llama.cpp [26] engine on the MUSICA supercomputer (Austria). Up to 80 NVIDIA H100 GPUs were employed concurrently during annotation. Each compute node with four H100 GPUs (80 GB VRAM each) handled batches of approximately 600,000 tokens per inference instance. Multi-node inference, which could allow for a higher context token limit, was not attempted as MUSICA is currently in its test phase; instead, parallelism was achieved by distributing text files across multiple nodes for simultaneous inference.
3.5 Evaluation Metrics for Hierarchical, non-Consistent Partitions
To quantitatively assess how well the inferred semantic trees align with existing labels in benchmark datasets, we employ a combination of external clustering validation measures: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI). For the evalution of these metrics we make use of their implementation in Scikit-Learn [29]. These metrics are applied independently at each tree layer, treating the clustering at that layer as a flat partition of the dataset.
ARI and NMI are widely used measures for comparing a predicted clustering against a ground-truth labeling. ARI evaluates agreement between two partitions based on pairwise co-assignment, adjusted for chance, while NMI measures the shared information between cluster assignments and class labels, normalized to account for differing numbers of clusters. In our setting, these metrics are computed per tree layer using only the documents assigned to clusters at that layer. This is necessary because the density-based procedure may leave a fraction of documents unclustered at higher density levels, leading to non-exhaustive and non-consistent partitions across layers. While ARI and NMI are not inherently designed for hierarchical or incomplete clusterings, they remain informative for assessing local alignment between inferred clusters and known labels where such overlap exists.